# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

### v4

@@ 3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will now see that hypothesis H, is also satisfied. From the equation (2), we have
$\psi_{w}(1,t)=t^{\Re(w)}\Big(1+(1-w)\int_{1}^{t}u^{-1-w}\,\{u\}\,du\Big),\quad\forall w\in\mathbb{C},\quad\forall t>1,$
@@
$\psi_{w}(1,t)$ $=t^{\Re(c)}\Big[1+J(t,\lfloor t\rfloor)+\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)\Big],$
$J_{w}(x,n)$ $:=(1-w)\int_{n}^{x}u^{-1-w}\,(u-n)\,du,$
with $\sum_{n=1}^{\lfloor t\rfloor-1}J_{w}(n+1,n)=0$ when $t\in[1,2)$. Use the integration by parts formula, we obtain
$J_{w}(x,n)$ $=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,\frac{d}{du}\Big(u^{-1}\,(u-n)\Big)du$
$=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,u^{-2}du.$
For every $\beta\in\mathbb{R}$ and $t\geq 2$ we have
$\Big|t^{\frac{1}{2}}J_{\frac{1}{2}+i\,\beta}(x,n)-J_{i\,\beta}(x,n)\Big|$ $\leq\int_{n}^{x}\Big(\big(t^{\frac{1}{2}}x^{\frac{1}{2}}-x\big)+(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)\,\Big)u^{-2}du$
Implies
$\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)$ $\geq t^{\frac{1}{2}}-1-\int_{\lfloor t\rfloor-1}^{t}(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)u^{-2}du$
$+\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big(n+1+u)\,u^{-2}du$
$-t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big((n+1)^{\frac{1}{2}}+u^{\frac{1}{2}}\big)u^{-2}du$
$\geq\frac{1}{3}\big(t^{\frac{1}{2}}-1\big)+\left(1-\frac{1}{t}\right),\qquad\forall t\geq 1.$
We recall that $\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)=0$ for all $t\in[1,2)$. We deduce that
$\Re\Big(\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)\Big)>0,\quad\forall t>1.$
Then the Hypothesis hypothesis H is satisfied. By definition of the function $\mu$ in the Definition 2, we have
@@
Since the Hypothesis hypothesis H is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.

### v6

@@ 3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will see in the following proposition that the hypothesis (H) is also satisfied
Proposition 7.
When $\eta(t)=\{t\}$ the rigidity hypothesis (H) is satisfied.
Proof.
Given the second item of Remark 6, we prove that : For all $\beta\in\mathbb{R}$ and $t\geq 1$ we have $\Re\big(f_{\beta}(t)\big)\geq 0$, where
$f_{\beta}(t):=\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du.$
We have
$f_{\beta}(t)=-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big]\,du,$
equivalent to
$f_{\beta}(t)=\int_{1}^{t}u^{-\frac{1}{2}}h_{\beta}(u)\,du=t^{-\frac{1}{2}}\int_{1}^{t}h_{\beta}(u)\,du+\frac{1}{2}\int_{1}^{t}u^{-\frac{3}{2}}\int_{1}^{u}h_{\beta}(v)\,dv\,du,$
where
$h_{\beta}(u):=-u^{-1-i\,\beta}\,\{u\}+u^{-\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big].$
Then it is sufficient to prove that $\int_{1}^{t}\Re\big(h_{\beta}(u)\big)\,du\geq 0$ for all $t\geq 1$. We have
$\int_{1}^{t}h_{\beta}(u)\,du$ $=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du$
$=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+2(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du-2(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du,$
We obtained
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du$
$-(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du.$
Since $\{u\}=u-\lfloor u\rfloor$, and since
$(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{1}{2}-i\,\beta}\,du-(1-i\beta)\int_{1}^{t}u^{-i\,\beta}\,\,du=0,$
then
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\lfloor u\rfloor\,du$
$+(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\lfloor u\rfloor\,du,$
@@
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1$
$-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\,du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-\frac{3}{2}-i\,\beta}\,du\Big]$
$+(1-i\beta)\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}u^{-1-i\,\beta}\ du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-1-i\,\beta}\,du\Big].$
Integrate, to obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-t^{\frac{1}{2}}\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}\lfloor t\rfloor\big(t^{-\frac{1}{2}-i\,\beta}-\lfloor t\rfloor^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\lfloor t\rfloor\big(t^{-i\,\beta}-\lfloor t\rfloor^{-i\,\beta}\big)$
$-\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-\frac{1}{2}-i\,\beta}-n^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-i\,\beta}-n^{-i\,\beta}\big).$
Using telescopic series, one obtains
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-\frac{1}{2}-i\,\beta}\Big)$
$+\frac{1-i\beta}{i\,\beta}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-i\,\beta}\Big),$
which is equivalent to,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1-i\beta}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
or,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)t^{1-i\,\beta}$
$-\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
Use the Abel summation formula to the last line, we obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}+t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}\,du$
$-\frac{1}{1-i\,\beta}-\int_{1}^{t}u^{-1-i\beta}\{u\}\,du.$
A direct estimate show that $\Re\big(f_{\beta}(t)\big)>0$, for all $t>1$. ∎
Then the rigidity hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
@@
Since the hypothesis (H) is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line. The results provides an answer to the Dynamical Conjecture formulated in [2].